Ticket sales for the Spring Fling totaled $1540.
Advance tickets were $3 each. At the door, tickets were $4 each. There were twice as many advance tickets sold as were sold at the door.
How many of each kind of ticket were sold?

Answer by TimothyLamb(4379)

x = advance tickets
y = door tickets
---
3x + 4y = 1540
x = 2y
---
put the system of linear equations into standard form
---
3x + 4y = 1540
x - 2y = 0
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = advance tickets = 308
y = door tickets = 154
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---